OpenStudy (anonymous):

Solve for x: 2(1-3x) - (x-5) > 1

OpenStudy (anonymous):

2^(3x-1)=5 taking log on both sides (3x-1)Ln2=Ln5 3x-1=Ln5/Ln2 But Ln5/Ln2=log(5) 2 [ ie read it as log 2 to base 5] theerefore 3x=log(5) 2 +1 simplifying x=(log(5)2+1)/3 truns out to be x=1.1

OpenStudy (anonymous):

I'm assuming that's just a 2, not a 2^(1-3x). It's not really clear. You want to simplify your equation down to it's answer. Multiply the 2 to the 1 and the -3x, and the - into (x-5) 2-6x-x+5>1 -7x+7>1 -7x>6 x<(-6/7)

OpenStudy (anonymous):

wait, if +7 is subtracted wouldn't the answer be positive 6/7

OpenStudy (anonymous):

wait yes you're right XD Sorry for the slip, you're right.